Friday, 21 August

Other Events

\underline{17:30 Urania} Klaus Becker, Klieve, Germany

From Stone–Sculpturing to Mathematics: Approximating the Inscribed Ball of Regular Polyhedra

\underline{18:00 Urania} Movie

N is a Number — a Portrait of Paul Erdős

\underline{19:30 Urania} Phil Davis, Brown University, Providence, USA

The Prospects of Mathematics in a Multimedia Civilization

\underline{20:00 Urania} VideoMath Festival

The “VideoMath Reel”, the 24 best contributions to the VideoMath competition

ICM’98 – Program  1